A car travels 450km in 4 hours. Calculate the average speed of a car in km/hr. (Round answer to one decimal place)
1
Expert's answer
2020-09-22T15:30:37-0400

We can find the average speed of the car from the formula:


"v = \dfrac{s}{t} = \dfrac{450 \ km}{4 \ h} = 112.5 \ \dfrac{km}{h}."

Answer:

"v = 112.5 \ \dfrac{km}{h}."